OpenStudy (anonymous):

in 1995 the life expectancy of males in a certian country was 64.2 yrs. In 2001 it was 66.5 yrs. Let E represent the life expectancy in Yr t and let t represent the number of yrs. since 1995. The linear function E (t) that fits the data is E(t)=? t+? round to nearest tenth Use the function to predict the life expectancy of males in 2004 E (9)=? round to the nearest tenth

OpenStudy (anonymous):

E (t)=0.5t+64.2 E(9)=0.5(9)=64.2=68.7 right?

OpenStudy (anonymous):

E(9) =0.5 (9)=64.2=68.7 rounded to the nearest tenth is 70